A chemist has two solutions H2SO4. One has a 40% concentration and the other one has a 25% concentration. How many liters of each solution must be mixed to obtain 106 liters of a 36% solution?
Answer by Maths68(1474)

Solution A
Amount = x liters
Concentration =40% = 0.40
=======================================================
Solution B
Amount =(106- x) liters
Concentration =25% = 0.25
=======================================================
Resultant Solution
Amount = 106 liters
Concentration = 36%=0.36
=======================================================
[Amount Solution A * Concentration A] + [Amount Solution B * Concentration of B] = Amount of Resultant * Concentration of resultant
========================================================
Put the value in above equation
(x*0.40)+((106-x)*0.25=106*0.36
0.40x+26.5-0.25x=38.16
0.15x=38.16-26.5
0.15x=11.66
Divide by 0.15 both sides
0.15x/0.15=11.66/0.15
x=77.73
=======================================
Amount of solution A = 77.73 liters
Amount of Sultion B = 106-77.73=28.27 liters